You note the following yield curve in The Wall Street Journal. According to the unbiased expectations theory, what is the one-year forward rate for the period beginning two years from today, 3f1? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

Maturity Yield
One day 2.24 %
One year 2.76
Two years 2.90
Three years 2.93
Answer #1

As per expectations theory, investing for 3 years at the 3-year rate should result in the same ending value as investing for 2 years at the 2-year rate, and reinvesting the proceeds after 2 years at the 1-year rate 2 years from now.

Let us say the 1-year rate 2 years from now is R. Then :

(1 + 2.93%)3 = (1 + 2.90%)2 * (1 + R)

R = (1.02933 / 1.0292) - 1

R = 2.99%

One-year forward rate for the period beginning two years from today 3f1 = 2.99%